OpenStudy (anonymous):

A treasure chest has a volume of x3 + 9x2 + 26x + 24 cm3 and the height is x + 4 cm. Find the polynomial that would represent the area of the bottom of the treasure chest? Explain your reasoning.

OpenStudy (anonymous):

volume = \[x ^{3} + 9x^{2} + 26x + 24\] height = x + 4 The bottom area of the chest would be length times width. Since: Volume = length x width x height \[\frac{length * width * height}{height} = length * width\]

OpenStudy (anonymous):

so just divide out x + 4?

OpenStudy (anonymous):

Yep. volume / height = area of the bottom